### v2

@@ 2.4 RL as a General Substrate for Decision-Making
A key feature of this phase is the emphasis on problem formulation over algorithmic novelty. In applications like RLHF [54], the choice of RL algorithm (e.g., PPO [99], GRPO [100, 101]) is secondary; the critical component is the learned reward model derived from human preferences. Similarly, in systems like AlphaGo [102], RL serves as one module within a broader architecture that includes imitation, self-play, and tree search. In these cases, RL acts as a fine-tuning layer that optimizes a well-defined objective.
@@ 3 RL for Image Generation
Reinforcement learning (RL) provides an effective framework for guiding image generation models toward desired outcomes by optimizing reward signals. Unlike purely supervised methods, RL excels at incorporating human preferences and maintaining semantic consistency between inputs (e.g., prompts or captions) and generated images. Current RL-based approaches for image generation fall into two main categories. The first includes policy-based methods, such as Proximal Policy Optimization (PPO) and recent variants like Group Relative Policy Optimization (GRPO), which optimize policies through gradient-based updates. The second is Direct Preference Optimization (DPO), which reformulates the problem as a preference classification task using ranked output pairs.
@@
DPOK[103] Policy Gradient May 2023 NeurIPS 2023
Promptist[104] Policy Gradient Dec 2022 NeurIPS 2023
DDPO[105] Policy Gradient May 2023 Arxiv
RLD[106] Policy Gradient May 2023 CVPR 2024
Diffusion-DPO[107] DPO Nov 2023 CVPR 2024
PRDP[108] DPO Feb 2024 CVPR 2024
DenseRewardT2I[109] DPO Feb 2024 ICML 2024
POSI[110] Policy Gradient Feb 2024 Arxiv
AGFSync[111] DPO Mar 2024 Arxiv
PAE[112] Policy Gradient Apr 2024 CVPR 2024
CurriculumDPO[113] DPO May 2024 CVPR 2025
HG-DPO[114] DPO May 2024 CVPR 2025
SPO[115] DPO Sat 2024 Arxiv
ReNO[116] DPO Jun 2024 NeurIPS 2024
DUO[117] DPO Jul 2024 NeurIPS 2024
RankDPO[118] DPO Oct 2024 Arxiv
PatchDPO[119] DPO Dec 2024 CVPR 2025
PPD[120] DPO Jan 2025 CVPR 2025
ImageGenerationCoT[121] DPO Jan 2025 Arxiv
CaPO[122] DPO Feb 2025 CVPR 2025
DesignDiffusion[123] DPO Mar 2025 CVPR 2025
LightGen[124] DPO Mar 2025 Arxiv
SimpleAR[125] GRPO Apr 2025 Arxiv
DanceGRPO[28] GRPO May 2025 Arxiv
Flow-GRPO[126] GRPO May 2025 Arxiv
T2I-R1[127] GRPO May 2025 Arxiv
ReasonGen-R1[128] GRPO May 2025 Arxiv
@@ 3.1 PPO-based Image Generation
One representative approach is Denoising Diffusion Policy Optimization (DDPO) [105], which formulates the denoising process in diffusion models as a multi-step MDP. This allows the model to directly optimize downstream generation objectives, rather than relying solely on likelihood-based training. Concurrently, DPOK [103] also advances this direction by framing the fine-tuning of text-to-image diffusion models as an online RL problem within a multi-step Markov Decision Process (MDP) framework. It optimizes feedback-trained rewards using policy gradient, critically incorporating Kullback–Leibler (KL) regularization to concurrently enhance both image quality and text-image alignment.
To promote output diversity, Wallace et al. [106] introduce reward functions based on Maximum Mean Discrepancy (MMD) and mutual information. These rewards evaluate the distributional coverage of generated images relative to reference sets. Each image receives marginal-utility-based feedback, encouraging the model to produce diverse outputs while maintaining sample quality.
Beyond generation-stage optimization, PPO-based methods have also been applied to prompt engineering. Promptist [104] proposes a two-stage pipeline that combines supervised fine-tuning (SFT) with PPO-based RL to automatically optimize user prompts. This improves alignment between user intent and the model’s internal preference signals. POSI [110] extends this idea to safe text-to-image generation by using a custom reward function that balances semantic alignment with safety constraints. PAE [112] further introduces dynamic prompt optimization, enabling fine-grained control over prompt weights and injection timing. By integrating SFT with online PPO, PAE optimizes for aesthetics, semantic fidelity, and user preference under a unified reward formulation.
@@ 3.2 DPO-based Image Generation
Direct Preference Optimization (DPO) has emerged as a highly effective method for aligning image generation models with human preferences. Early work applied DPO to standard generative architectures. For example, Diffusion-DPO [107] and PRDP [108] align diffusion models with preference data through post-training optimization, improving both output quality and controllability.
DPO has also been extended to specific sub-tasks. HG-DPO [114] addresses structural and pose inaccuracies in human image generation by using high-quality real images as reference preferences. RPO introduces a harmonic reward and early stopping to improve theme-guided generation while reducing overfitting. PatchDPO [119] incorporates patch-level feedback for personalized generation, enhancing local consistency and detail fidelity. Similarly, DesignDiffusion [123] applies DPO to produce high-quality design-oriented images from text prompts.
In addition to task-specific refinements, several works apply DPO to broader challenges in text-to-image generation. ReNO [116] improves compositional detail by optimizing initial noise during inference, helping models generalize beyond fine-tuning. DUO [117] reframes content safety as a preference optimization task, aiming to suppress unsafe outputs. SPO [115] focuses on aesthetic alignment by optimizing for visual appeal based on preference signals.
Some works refocus on enhancing the DPO framework itself. AGFSync [111] reduces dependence on manual labels by generating preference annotations via AI models. RankDPO [118] and CaPO [122] propose scalable modeling techniques to address the limitations of aging or costly human preference data. Dense Reward [109] introduces dense reward signals to capture fine-grained human judgments. Curriculum-DPO [113] adopts a curriculum learning approach, improving sample efficiency by gradually increasing task difficulty. PPD [120] extends DPO to model individual user preferences, addressing the limitations of population-level supervision.
Finally, DPO has also been applied to autoregressive models. LightGen [124] uses post-training DPO to refine generation quality. Image-Generation-CoT [121] combines DPO with Chain-of-Thought (CoT) prompting, enabling preference alignment in sequential generation tasks. These efforts reflect the growing utility and flexibility of DPO for aligning image generation models with nuanced, user-centered objectives.
@@ 3.3 GRPO-based Image Generation
Group Relative Policy Optimization (GRPO) [100] is a recent extension of policy optimization that offers improved generalization and training stability. As a variant of PPO, GRPO has gained traction in visual content generation and has been applied across diffusion- and flow-based pipelines.
DanceGRPO [28] proposes a unified GRPO framework that supports diverse generative paradigms, including diffusion models and rectified flows. It is applicable across multiple tasks such as text-to-image, text-to-video, and image-to-video generation. DanceGRPO supports various backbone models (e.g., Stable Diffusion, FLUX) and reward types (e.g., aesthetics, alignment, motion), demonstrating the flexibility of GRPO in unifying heterogeneous tasks and architectures under a single policy-based training scheme.
Flow-GRPO [126] extends GRPO to flow-based generative models by reformulating generation as a stochastic differential equation (SDE). This allows for effective RL-based exploration. It also introduces a denoising-step shrinking strategy to improve both sample efficiency and output fidelity.
GRPO has also been applied in autoregressive generation. T2I-R1 [127] introduces BiCoT-GRPO, a joint optimization framework that aligns both semantic-level and token-level Chain-of-Thought (CoT) reasoning within a single GRPO loop. SimpleAR [125] shows that GRPO can significantly enhance prompt-image alignment and visual quality when combined with supervised fine-tuning. ReasonGen-R1 [128] further integrates CoT reasoning into GRPO training, aiming to improve logical coherence across multi-step generation processes.
@@ 4 RL for Video Generation
AdaDiff [129] Policy Gradient Nov 2023 AAAI 2025
RLAVE [130] Actor-Critic Nov 2024 MM 2023
InstructVideo [131] Reward Fine-tuning Dec 2023 CVPR 2024
VADER [132] Reward Gradient Jul 2024 arXiv
SePPO [133] DPO Oct 2024 arXiv
VideoAgent [134] Online Policy Fine-tuning Oct 2024 ECCV 2024
RL-V2V-GAN [135] Policy Gradient Oct 2024 arXiv
E-Motion [136] PPO Oct 2024 NeurIPS 2024
Free2Guide [137] Path Integral Control Nov 2024 arXiv
IDOI [138] RWR, DPO Dec 2024 arXiv
FLIP [139] Actor-Critic Dec 2024 ICLR2025
VideoDPO [140] DPO Dec 2024 CVPR 2025
OnlineVPO [141] DPO Dec 2024 arXiv
VisionReward [142] DPO Dec 2024 arXiv
HuViDPO [143] DPO Feb 2025 arXiv
IPO [144] DPO Feb 2025 arXiv
HALO [145] DPO Feb 2025 arXiv
AAVG [146] DPO Apr 2025 arXiv
SkyReelsV2 [147] DPO Apr 2025 arXiv
Phys-AR [148] GRPO Apr 2025 arXiv
DanceGRPO [28] GRPO May 2025 arXiv
Diffusion-NPO [149] NPO May 2025 ICLR 2025
RLVR-World [150] GRPO May 2025 arXiv
InfLVG [151] GRPO May 2025 arXiv
RDPO [7] DPO Jun 2025 arXiv
DenseDPO [152] DPO Jun 2025 arXiv
EchoMimicV3 [153] DPO Jul 2025 arXiv
@@ 4.1 Improving Diffusion Sampling Efficiency
One practical application of RL in video generation is optimizing the diffusion sampling process to reduce computational cost without sacrificing output quality. AdaDiff [129] proposes an adaptive step selection method that accelerates sampling by learning to adjust denoising step sizes dynamically. The process is framed as an RL problem, where the state is the current noisy video frame and timestep, the action is the size of the next denoising step, and the reward reflects both generation speed and final output quality. A policy network is trained to select step sizes that balance efficiency and fidelity. Larger steps are chosen when coarse updates suffice, while smaller steps are used when fine details are required. This adaptive scheduling enables faster, high-quality video generation with fewer denoising iterations.
@@ 4.2 Sequential Control: Planning and Editing
In video planning, RL guides generation toward task-specific goals. FLIP [139] formulates instructional video generation as a clip selection problem. It adopts an actor-critic framework, where the policy selects video segments to fulfill a textual instruction. A vision-language model provides dense feedback as the reward signal, allowing the system to learn semantically aligned and visually coherent video plans. VideoAgent [134] approaches planning from an embodied AI perspective. It uses demonstration-based videos to bootstrap a visual planner, which is later refined by executing rollouts in real environments. Although RL underpins the learning loop, the method avoids explicit policy gradients and instead uses reward-guided data selection.
More than planning, RL is also applied in video editing and transformation. RL-V2V-GAN [135] performs unsupervised video-to-video translation—for example, transforming summer scenes into winter. It incorporates policy gradients within a GAN framework, rewarding outputs with consistent temporal structure and style fidelity. This allows the model to learn coherent transformations without paired training data. E-Motion [136] focuses on future motion prediction using event-camera inputs. It frames the reverse process of diffusion as a MDP and applies PPO to fine-tune motion trajectories using perceptual rewards such as FVD and SSIM.
RL has also been extended to general-purpose video editing. RLAVE [130] proposes a framework where the actor selects editing operations—such as cuts, transitions, or clip selection—and the critic evaluates narrative coherence, pacing, and aesthetics using a vision-language model. The system is rewarded for producing stylistically consistent and engaging video edits, bridging the gap between automatic synthesis and human-guided post-production.

### v3

@@ 2.4 RL as a General Substrate for Decision-Making
A key feature of this phase is the emphasis on problem formulation over algorithmic novelty. In applications like RLHF [ouyang2022training], the choice of RL algorithm (e.g., PPO [engstrom2019implementation], GRPO [shao2024deepseekmath, chen2025seed]) is secondary; the critical component is the learned reward model derived from human preferences. Similarly, in systems like AlphaGo [granter2017alphago], RL serves as one module within a broader architecture that includes imitation, self-play, and tree search. In these cases, RL acts as a fine-tuning layer that optimizes a well-defined objective.
@@
Figure 2: RL-style post-training pipelines for diffusion models. Given a conditioning input $q$, the diffusion policy $\pi_{\theta}$ generates samples via iterative denoising. DPO optimizes pairwise preferences using denoiser outputs (e.g., $\epsilon$) relative to a frozen reference policy. PPO treats the denoising path $o$ as a policy rollout and updates the policy using advantage estimates. GRPO extends this formulation by computing relative advantages from a group of diffusion rollouts. All methods employ KL regularization to the reference policy to stabilize training.
@@ 3 RL for Image Generation
Reinforcement learning (RL) offers a principled framework to steer generative models—particularly diffusion models—toward objectives that are difficult to optimize with standard supervised losses. These include non-differentiable metrics such as aesthetic quality, precise semantic alignment, and compliance with safety guidelines. RL addresses these challenges by reframing the generative process as a sequential decision-making task, where reward signals guide the model to satisfy complex, often preference-driven goals. As a result, RL provides an effective mechanism for guiding image generation models toward desired outcomes by optimizing such reward signals. Unlike purely supervised methods, RL excels at incorporating human preferences and maintaining semantic consistency between inputs (e.g., prompts or captions) and generated images. Current RL-based approaches for image generation fall into two main categories. The first includes policy-based methods, such as Proximal Policy Optimization (PPO) and recent variants like Group Relative Policy Optimization (GRPO), which optimize policies through gradient-based updates. The second is Direct Preference Optimization (DPO), which reformulates the problem as a preference classification task using ranked output pairs. A chronological overview of representative RL methods for image generation is provided in Table 1.
@@
DPOK[fan2023dpok] Policy Gradient May 2023 NeurIPS 2023
Promptist[hao2023optimizing] Policy Gradient Dec 2022 NeurIPS 2023
DDPO[black2023training] Policy Gradient May 2023 Arxiv
RLD[miao2024training] Policy Gradient May 2023 CVPR 2024
Diffusion-DPO[wallace2024diffusion] DPO Nov 2023 CVPR 2024
PRDP[deng2024prdp] DPO Feb 2024 CVPR 2024
DenseRewardT2I[yang2024dense] DPO Feb 2024 ICML 2024
POSI[wu2024universal] Policy Gradient Feb 2024 Arxiv
AGFSync[an2025agfsync] DPO Mar 2024 Arxiv
PAE[mo2024dynamic] Policy Gradient Apr 2024 CVPR 2024
CurriculumDPO[croitoru2025curriculum] DPO May 2024 CVPR 2025
HG-DPO[na2025boost] DPO May 2024 CVPR 2025
SPO[liang2024step] DPO Sat 2024 Arxiv
ReNO[eyring2024reno] DPO Jun 2024 NeurIPS 2024
DUO[park2024direct] DPO Jul 2024 NeurIPS 2024
RankDPO[karthik2024scalable] DPO Oct 2024 Arxiv
PatchDPO[huang2025patchdpo] DPO Dec 2024 CVPR 2025
PPD[dang2025personalized] DPO Jan 2025 CVPR 2025
ImageGenerationCoT[guo2025can] DPO Jan 2025 Arxiv
CaPO[lee2025calibrated] DPO Feb 2025 CVPR 2025
DesignDiffusion[wang2025designdiffusion] DPO Mar 2025 CVPR 2025
LightGen[wu2025lightgen] DPO Mar 2025 Arxiv
SimpleAR[wang2025simplear] GRPO Apr 2025 Arxiv
DanceGRPO[dancegrpo] GRPO May 2025 Arxiv
Flow-GRPO[liu2025flow] GRPO May 2025 Arxiv
T2I-R1[jiang2025t2i] GRPO May 2025 Arxiv
ReasonGen-R1[zhang2025reasongen] GRPO May 2025 Arxiv
@@ 3.1 PPO-based Image Generation
One representative approach is Denoising Diffusion Policy Optimization (DDPO) [black2023training], which formulates the denoising process in diffusion models as a multi-step MDP. This allows the model to directly optimize downstream generation objectives, rather than relying solely on likelihood-based training. Concurrently, DPOK [fan2023dpok] also advances this direction by framing the fine-tuning of text-to-image diffusion models as an online RL problem within a multi-step Markov Decision Process (MDP) framework. It optimizes feedback-trained rewards using policy gradient, critically incorporating Kullback–Leibler (KL) regularization to concurrently enhance both image quality and text-image alignment. To address the sparse-reward challenge in such policy-gradient approaches, B²-DiffuRL [hu2025towards] introduces a backward progressive training scheme with branch-based sampling, which densifies the learning signal and enables more precise policy updates, leading to better alignment-diversity trade-offs. Separately, LOOP [gupta2025simple] enhances training stability and efficiency by employing a leave-one-out baseline across multiple trajectories, effectively reducing variance and memory overhead while preserving the clipping mechanism of PPO.
To promote output diversity, Wallace et al. [miao2024training] introduce reward functions based on Maximum Mean Discrepancy (MMD) and mutual information. These rewards evaluate the distributional coverage of generated images relative to reference sets. Each image receives marginal-utility-based feedback, encouraging the model to produce diverse outputs while maintaining sample quality.
Beyond generation-stage optimization, PPO-based methods have also been applied to prompt engineering. Promptist [hao2023optimizing] proposes a two-stage pipeline that combines supervised fine-tuning (SFT) with PPO-based RL to automatically optimize user prompts. This improves alignment between user intent and the model’s internal preference signals. POSI [wu2024universal] extends this idea to safe text-to-image generation by using a custom reward function that balances semantic alignment with safety constraints. PAE [mo2024dynamic] further introduces dynamic prompt optimization, enabling fine-grained control over prompt weights and injection timing. By integrating SFT with online PPO, PAE optimizes for aesthetics, semantic fidelity, and user preference under a unified reward formulation.
@@ 3.2 DPO-based Image Generation
Direct Preference Optimization (DPO) has emerged as a highly effective method for aligning image generation models with human preferences. Early work applied DPO to standard generative architectures. For example, Diffusion-DPO [wallace2024diffusion] and PRDP [deng2024prdp] align diffusion models with preference data through post-training optimization, improving both output quality and controllability.
DPO has also been extended to specific sub-tasks. HG-DPO [na2025boost] addresses structural and pose inaccuracies in human image generation by using high-quality real images as reference preferences. RPO introduces a harmonic reward and early stopping to improve theme-guided generation while reducing overfitting. PatchDPO [huang2025patchdpo] incorporates patch-level feedback for personalized generation, enhancing local consistency and detail fidelity. Similarly, DesignDiffusion [wang2025designdiffusion] applies DPO to produce high-quality design-oriented images from text prompts.
In addition to task-specific refinements, several works apply DPO to broader challenges in text-to-image generation. ReNO [eyring2024reno] improves compositional detail by optimizing initial noise during inference, helping models generalize beyond fine-tuning. DUO [park2024direct] reframes content safety as a preference optimization task, aiming to suppress unsafe outputs. SPO [liang2024step] focuses on aesthetic alignment by optimizing for visual appeal based on preference signals.
Some works refocus on enhancing the DPO framework itself. AGFSync [an2025agfsync] reduces dependence on manual labels by generating preference annotations via AI models. RankDPO [karthik2024scalable] and CaPO [lee2025calibrated] propose scalable modeling techniques to address the limitations of aging or costly human preference data. Dense Reward [yang2024dense] introduces dense reward signals to capture fine-grained human judgments. Curriculum-DPO [croitoru2025curriculum] adopts a curriculum learning approach, improving sample efficiency by gradually increasing task difficulty. PPD [dang2025personalized] extends DPO to model individual user preferences, addressing the limitations of population-level supervision.
Finally, DPO has also been applied to autoregressive models. LightGen [wu2025lightgen] uses post-training DPO to refine generation quality. Image-Generation-CoT [guo2025can] combines DPO with Chain-of-Thought (CoT) prompting, enabling preference alignment in sequential generation tasks. These efforts reflect the growing utility and flexibility of DPO for aligning image generation models with nuanced, user-centered objectives.
@@ 3.3 GRPO-based Image Generation
Group Relative Policy Optimization (GRPO) [shao2024deepseekmath] is a recent extension of policy optimization that offers improved generalization and training stability. As a variant of PPO, GRPO has gained traction in visual content generation and has been applied across diffusion- and flow-based pipelines.
DanceGRPO [dancegrpo] proposes a unified GRPO framework that supports diverse generative paradigms, including diffusion models and rectified flows. It is applicable across multiple tasks such as text-to-image, text-to-video, and image-to-video generation. DanceGRPO supports various backbone models (e.g., Stable Diffusion, FLUX) and reward types (e.g., aesthetics, alignment, motion), demonstrating the flexibility of GRPO in unifying heterogeneous tasks and architectures under a single policy-based training scheme.
Flow-GRPO [liu2025flow] extends GRPO to flow-based generative models by reformulating generation as a stochastic differential equation (SDE). This allows for effective RL-based exploration. It also introduces a denoising-step shrinking strategy to improve both sample efficiency and output fidelity. Further extending its application, TALK2MOVE [tan2026talk2move] applies Flow-GRPO to instruction-based geometric image editing, utilizing a spatially-grounded reward function based on segmentation and depth estimates to achieve precise object manipulation without paired training data.
The application of GRPO also extends to autoregressive generation paradigms. T2I-R1 [jiang2025t2i] introduces BiCoT-GRPO, a joint optimization framework that aligns both semantic-level and token-level Chain-of-Thought (CoT) reasoning within a single GRPO loop. SimpleAR [wang2025simplear] demonstrates that integrating GRPO with supervised fine-tuning can significantly enhance prompt-image alignment and visual quality. Building on this, ReasonGen-R1 [zhang2025reasongen] further incorporates CoT reasoning into the GRPO training process to improve logical coherence in multi-step generation.
@@ 4 RL for Video Generation
AdaDiff [zhang2025adadiff] Policy Gradient Nov 2023 AAAI 2025
RLAVE [hu2023reinforcement] Actor-Critic Nov 2024 MM 2023
InstructVideo [instructvideo] Reward Fine-tuning Dec 2023 CVPR 2024
VADER [vader] Reward Gradient Jul 2024 arXiv
SePPO [sspo] DPO Oct 2024 arXiv
VideoAgent [soni2025videoagent] Online Policy Fine-tuning Oct 2024 ECCV 2024
RL-V2V-GAN [v2vganrl] Policy Gradient Oct 2024 arXiv
E-Motion [emotion] PPO Oct 2024 NeurIPS 2024
Free2Guide [kim2024free] Path Integral Control Nov 2024 arXiv
IDOI [furuta2024improving] RWR, DPO Dec 2024 arXiv
FLIP [gao2024flip] Actor-Critic Dec 2024 ICLR2025
VideoDPO [videodpo] DPO Dec 2024 CVPR 2025
OnlineVPO [zhang2024onlinevpo] DPO Dec 2024 arXiv
VisionReward [visionreward] DPO Dec 2024 arXiv
HuViDPO [jiang2025huvidpo] DPO Feb 2025 arXiv
IPO [yang2025ipo] DPO Feb 2025 arXiv
HALO [wang2025harness] DPO Feb 2025 arXiv
AAVG [zhu2025aligning] DPO Apr 2025 arXiv
SkyReelsV2 [skyreels] DPO Apr 2025 arXiv
Phys-AR [physar] GRPO Apr 2025 arXiv
DanceGRPO [dancegrpo] GRPO May 2025 arXiv
Diffusion-NPO [wang2025diffusion] NPO May 2025 ICLR 2025
RLVR-World [rlvrworld] GRPO May 2025 arXiv
InfLVG [inflvg] GRPO May 2025 arXiv
RDPO [rdpo] DPO Jun 2025 arXiv
DenseDPO [densedpo] DPO Jun 2025 arXiv
EchoMimicV3 [echomimicv3] DPO Jul 2025 arXiv
@@ 4.1 Improving Diffusion Sampling Efficiency
One practical application of RL in video generation is optimizing the diffusion sampling process to reduce computational cost without sacrificing output quality. AdaDiff [zhang2025adadiff] proposes an adaptive step selection method that accelerates sampling by learning to adjust denoising step sizes dynamically. The process is framed as an RL problem, where the state is the current noisy video frame and timestep, the action is the size of the next denoising step, and the reward reflects both generation speed and final output quality. A policy network is trained to select step sizes that balance efficiency and fidelity. Larger steps are chosen when coarse updates suffice, while smaller steps are used when fine details are required. This adaptive scheduling enables faster, high-quality video generation with fewer denoising iterations.
@@ 4.2 Sequential Control: Planning and Editing
In video planning, RL guides generation toward task-specific goals. FLIP [gao2024flip] formulates instructional video generation as a clip selection problem. It adopts an actor-critic framework, where the policy selects video segments to fulfill a textual instruction. A vision-language model provides dense feedback as the reward signal, allowing the system to learn semantically aligned and visually coherent video plans. VideoAgent [soni2025videoagent] approaches planning from an embodied AI perspective. It uses demonstration-based videos to bootstrap a visual planner, which is later refined by executing rollouts in real environments. Although RL underpins the learning loop, the method avoids explicit policy gradients and instead uses reward-guided data selection.
More than planning, RL is also applied in video editing and transformation. RL-V2V-GAN [v2vganrl] performs unsupervised video-to-video translation—for example, transforming summer scenes into winter. It incorporates policy gradients within a GAN framework, rewarding outputs with consistent temporal structure and style fidelity. This allows the model to learn coherent transformations without paired training data. E-Motion [emotion] focuses on future motion prediction using event-camera inputs. It frames the reverse process of diffusion as a MDP and applies PPO to fine-tune motion trajectories using perceptual rewards such as FVD and SSIM.
RL has also been extended to general-purpose video editing. RLAVE [hu2023reinforcement] proposes a framework where the actor selects editing operations—such as cuts, transitions, or clip selection—and the critic evaluates narrative coherence, pacing, and aesthetics using a vision-language model. The system is rewarded for producing stylistically consistent and engaging video edits, bridging the gap between automatic synthesis and human-guided post-production.
Table 3: Commonly Used Reward Models in Visual Content Generation
Model Name Modality Backbone Typical Use
LAION Aesthetic Predictor [laion_aesthetic] Image ViT Aesthetic quality reward
CLIPScore [radford2021learning] Text-Image ViT Cross-modal similarity reward
BLIPScore [li2022blip] Vision-Language ViT Semantic consistency reward
VideoCLIP-XL [wang2024videoclip] Text-Video ViT Text-video alignment reward
PickScore [kirstain2023pick] Text-Image ViT Pairwise preference reward
HPSv2 [wu2023human] Text-Image ViT Human preference reward
UnifiedReward [wang2025unified] Image + Video LLM Unified generation reward
IPO [yang2025ipo] Text-Video LLM Iterative T2V alignment reward
LiFT [wang2024lift] Text-Video LLM Human-feedback alignment reward
VideoScore [he2024videoscore] Video LLM Video quality assessment reward
VisionReward [visionreward] Image + Video LLM Fine-grained VQA-based reward
VideoAlign[liu2025improving] Text-Video LLM Multidimensional T2V quality reward
AnimeReward [zhu2025aligning] Image-to-Video LLM Anime video preference reward